Mr Brown gave a test with 30 questions worth 100 points. If some of the questions were worth three points and the rest were worth 5 points, How many questions of each value did he have?
Answer by lwsshak3(11628)

Mr Brown gave a test with 30 questions worth 100 points. If some of the questions were worth three points and the rest were worth 5 points, How many questions of each value did he have?
***
let x=number of 5-point questions
30-x=number of 3-point questions
..
5x+3(30-x)=100
5x+90-3x=100
2x=10
x=5
30-x=25
number of 5-point questions=5
number of 3-point questions=25